[Brahms-l] Shift report 20050422 00:00-08:00

From: <apache@brahms-web.brahms.bnl.gov>
Date: Fri Apr 22 2005 - 08:01:46 EDT
Shifter leader was kh
Shifters were 


No beam the entire night.  There were Access control problems, then Linac problems.  It looks as if it might be getting close now at shift end.

_______________________________________________
Brahms-l mailing list
Brahms-l@lists.bnl.gov
http://lists.bnl.gov/mailman/listinfo/brahms-l
Received on Fri Apr 22 08:02:05 2005

This archive was generated by hypermail 2.1.8 : Fri Apr 22 2005 - 08:02:20 EDT